OpenStudy (anonymous):

The base of a cube is 36 cm2. What is the volume of the cube?

OpenStudy (anonymous):

please show your work

OpenStudy (cwrw238):

if base area is 36cm2 then each side = 6 cm long and volume of cube = 6^3 or volume of cube = base times height = 36* 6

OpenStudy (anonymous):

Thank You

OpenStudy (cwrw238):

yw